# Exploring Dark Knowledge under Various Teacher Capacities and Addressing Capacity Mismatch

### v1

@@ VI-A Performance Comparisons
We compare with SOTA methods and list the results on CIFAR-100, TinyImageNet, CUB, and Dogs in Tab. II. The compared methods include ESKD [32], TAKD [33], SCKD [34], and ATS [31]. NoKD trains students without the teacher’s supervision. ST-KD trains students under the guidance of a smaller teacher. KD trains students under the guidance of the larger teacher. The larger teachers are ResNet110, WRN50-2, RNX101-32-8d, and RNX101-32-8d for the four datasets, while the smaller teachers are ResNet20, WRN26-2, RNX50-32-4d, and RNX50-32-4d, respectively. The values in “()” display the test accuracy of the large teachers. The last column of the table shows the average performance of corresponding rows. The last three rows present the performances of our proposed algorithms. These methods could improve the performances of students when taught by a larger network and surpass that of KD and ST-KD, which verifies that enhancing the variance of non-ground-truth classes could indeed make larger networks teach well again. The proposed ISATS could achieve the SOTA results when compared with ATS. It is easy to understand that ISATS searches for optimal temperatures in an instance-specific manner, which could enhance the variance of non-ground-truth class probabilities more effectively. RegT belongs to the pre-processing method that enhances the variance of non-ground-truth classes during the phase of training teachers, which also presents good performances.
@@
We also verify the ensemble performances after distillation because ResKD [35, 36] improves the students’ performances by introducing the residual student and taking the two residual students’ ensemble. Although ResKD surpasses the performance of ATS, this comparison is not fair. Hence, we also provide ensemble performances of two separately trained students under different initialization and training seeds. For a fair comparison, we also test the performances of our methods by repeating the corresponding algorithms two times and making predictions via the ensemble of the obtained two student networks. The results are listed in Tab. III. Clearly, our proposed ISATS could also achieve better performances when compared with ATS and other methods.
@@
Figure 8: Distillation results via TS, ATS, and ISATS on CIFAR-100. The x-axis of each figure shows teachers with various capacities.
VI-B Comparing ISATS with TS and ATS
@@
We then especially focus on the comparison of ISATS with TS and ATS. Specifically, we follow the experimental settings in [31] and additionally plot the performance curves when utilizing our proposed ISATS. Fig. 8 shows several pairs of KD experimental studies, which include the combinations of three series of teacher networks (i.e., ResNet, WideResNet, and ResNeXt) and three student networks (i.e., VGG8, ShuffleNetV1, and MobileNetV2). The utilized dataset is CIFAR-100. Each plot shows three curves, including TS (i.e., KD), ATS (i.e., KD+TS), and ISATS (i.e., KD+ATS). The x-axis represents the increasing capacity of teachers. The curves of TS imply that the student’s performance taught by teacher networks becomes worse when the teacher capacity increases, i.e., the capacity mismatch phenomenon. ATS could mitigate this phenomenon and make larger teachers teach well or better again. Our proposed ISATS has the same effect and could surpass ATS in most cases.
@@
Figure 9: The teacher’s label under different levels of confidence. The dark knowledge provided by the teacher with proper confidence is preferred.
VII Discussion
@@
  • Network Capacity. This paper mainly explores this factor. An extremely smaller network tends to provide uniform predictions, while an extremely larger network provides “one-hot” ones.
  • Training Process. With the training process of a network, it becomes more and more confident. In the beginning, it is ignorant and provides uniform predictions. After enough training steps, it may output “one-hot” vectors.
  • Temperature Factor. Given a trained network, we could also adjust the temperature and obtain different types of probability vectors. If we take a very high temperature, the distribution becomes uniform. In contrast, a lower temperature leads to “one-hot” results.
As shown in Fig. 9, a proper confidence level could generate a rational probability distribution that facilitates the KD process. For example, traditional KD methods do not use a too-small or too-large temperature [9]. Some works also point out that intermediate model checkpoints could be better teachers than the fully converged model [51]. That is, a proper training level of teachers also matters a lot in the KD process. Our work points out that the teacher with high capacity does not perform well in KD and proposes several solutions to adjust their over-confident outputs.
VIII Conclusion
This paper studies the dark knowledge in KD under various teacher capacities. Two observations are first presented, i.e., the probability vectors provided by larger teachers are less distinct between non-ground-truth classes, while the relative probability values seem consistent between different teachers. Abundant experimental studies are provided to explain these two observations. The less varied non-ground-truth class probabilities make the student hardly grasp the absolute affinities of non-ground-truth classes to the target class, leading to the interesting “capacity mismatch” phenomenon in KD. Multiple effective and simple solutions are proposed to solve this problem. Finally, a unified perspective about dark knowledge under various confidence levels is provided for future studies.

### v2

@@ 6.1 Performance Comparisons
We compare with SOTA methods and list the results on CIFAR-100, TinyImageNet, CUB, and Dogs in Tab. 2. The compared methods include ESKD [32], TAKD [33], SCKD [34], and ATS [31]. NoKD trains students without the teacher’s supervision. ST-KD trains students under the guidance of a smaller teacher. KD trains students under the guidance of the larger teacher. The larger teachers are ResNet110, WRN50-2, RNX101-32-8d, and RNX101-32-8d for the four datasets, while the smaller teachers are ResNet20, WRN26-2, RNX50-32-4d, and RNX50-32-4d, respectively. The values in “()” display the test accuracy of the large teachers. The last column of the table shows the average performance of corresponding rows. The last three rows present the performances of our proposed algorithms. These methods could improve the performances of students when taught by a larger network and surpass that of KD and ST-KD, which verifies that enhancing the variance of non-ground-truth classes could indeed make larger networks teach well again. Although dataset bias exists in a few circumstances, the proposed ISATS could achieve the SOTA results when compared with ATS. It is easy to understand that ISATS searches for optimal temperatures in an instance-specific manner, which could enhance the variance of non-ground-truth class probabilities more effectively. The two methods presented in Sections 5.1 and 5.2 are also meaningful. RegT belongs to the pre-processing method that enhances the variance of non-ground-truth classes during the phase of training teachers, which also presents good performances. RegT does not directly affect the teacher model’s error class variance like ATS, so its performance is not as good as ATS. However, together with ESKD [32], it indicates that a more suitable teacher is benefical to KD performance improvement and provides a simple and effective method for teacher regularization. Similarly, the performance improvement of FGCR is less pronounced than ATS, but it indirectly supports the rationale for moderate LS. The experimental data for these two methods, from the perspective of wrong-class variance that we introduced, demonstrate the validity of previous works that tried to address capacity mismatch.
@@
We also verify the ensemble performances after distillation because ResKD [35, 36] improves the students’ performances by introducing the residual student and taking the two residual students’ ensemble. Although ResKD surpasses the performance of ATS, this comparison is not fair. Hence, we also provide ensemble performances of two separately trained students under different initialization and training seeds. For a fair comparison, we also test the performances of our methods by repeating the corresponding algorithms two times and making predictions via the ensemble of the obtained two student networks. The results are listed in Tab. 3. Clearly, our proposed ISATS could also achieve better performances when compared with ATS and other methods.
@@
Figure 8: Distillation results via TS, ATS, and ISATS on CIFAR-100. (a)ResNet teachers;(b)WideResNet teachers;(c)ResNeXt teachers
6.2 Comparing ISATS with TS and ATS
@@
We then especially focus on the comparison of ISATS with TS and ATS. Specifically, we follow the experimental settings in [31] and additionally plot the performance curves when utilizing our proposed ISATS. Fig. 8 shows several pairs of KD experimental studies, which include the combinations of three series of teacher networks (i.e., ResNet, WideResNet, and ResNeXt) and three student networks (i.e., VGG8, ShuffleNetV1, and MobileNetV2). The utilized dataset is CIFAR-100. Each plot shows three curves, including TS (i.e., KD), ATS (i.e., KD+ATS), and ISATS (i.e., KD+ISATS). The x-axis represents the increasing capacity of teachers. The curves of TS imply that the student’s performance taught by teacher networks becomes worse when the teacher capacity increases, i.e., the capacity mismatch phenomenon. ATS could mitigate this phenomenon and make larger teachers teach well or better again. Our proposed ISATS has the same effect and could surpass ATS in most cases.
Fig. 8 further demonstrates that ISATS can achieve superior KD performance with a given teacher model size, avoiding the substantial effort required to find an appropriate teacher model in the situation of TS. Notably, ISATS outperforms ATS in distillation performance, particularly when the teacher model is large.
@@
Figure 9: The teacher’s label under different levels of confidence. The dark knowledge provided by the teacher with proper confidence is preferred for reaching the highest derived variance (variance of probabilities of non-ground-truth classes).
7 Discussion
@@
  • Network Capacity. This paper mainly explores this factor. An extremely smaller network tends to provide uniform predictions, while an extremely larger network provides “one-hot" ones.
  • Training Process. With the training process of a network, it becomes more and more confident. In the beginning, it is ignorant and provides uniform predictions. After enough training steps, it may output “one-hot" vectors. That is, a proper training level of teachers also matters a lot in the KD process. Some works have pointed out that intermediate model checkpoints could be better teachers than the fully converged model [53].
  • Temperature Factor. Given a trained network, we could also adjust the temperature and obtain different types of probability vectors. If we take a very high temperature, the distribution becomes uniform. In contrast, a lower temperature leads to “one-hot" results. Consequently, traditional KD methods do not use a too-small or too-large temperature [9].
As shown in Fig. 9, a proper confidence level could generate a rational probability distribution that facilitates the KD process. Our work points out why the teacher with high capacity does not perform well in KD and proposes several solutions to adjust their over-confident outputs.
8 Conclusion
This paper studies the dark knowledge in KD under various teacher capacities. Two observations are first presented, i.e., the probability vectors provided by larger teachers are less distinct among non-ground-truth classes, while the relative probability values seem consistent among different teachers. Abundant experimental studies are provided to explain these two observations. The less varied non-ground-truth class probabilities make the student hardly grasp the absolute affinities of non-ground-truth classes to the target class, leading to the interesting “capacity mismatch” phenomenon in KD. Multiple effective and simple solutions are proposed to solve this problem. Finally, a unified perspective about dark knowledge under various confidence levels is provided for future studies.
[…]
This work is supported by National Science and Technology Major Project (Grant No. 2022ZD0114805).